OpenStudy (anonymous):

Sophia has 8 books in her locker. All the books are either school books or personal books. She has three times as many school books than personal books. How many school books does Sophia have in her locker ?

OpenStudy (anonymous):

a-3b=8

OpenStudy (anonymous):

a+3b=8 oops

OpenStudy (anonymous):

a=8-3b

OpenStudy (anonymous):

well that works but is incorrect since you have 2 variable

OpenStudy (anonymous):

all the books are not the same

OpenStudy (anonymous):

hence a+3(b)=8

OpenStudy (anonymous):

since she has three times as many school books